Suppose junk emails delivered to one's inbox follow a poisson distribution, with a constante average rate of 3 emails per hour. What is the probability that one junk email is received within half an hour after creating an email account?
Answer #1

Ans:

Given that

Average rate=3 emails per hr

So,

There will be on average 3*0.5=1.5 emails in half an hr

P(x=1)=e-1.5*(1.51/1!)=0.3347